# Saddle Hierarchy in Dense Associative Memory

### v1

@@ Abstract
Dense associative memory (DAM) models have been attracting renewed attention since they were shown to be robust to adversarial examples and closely related to state-of-the-art machine learning paradigms, such as the attention mechanisms in transformers and generative diffusion models. We study a DAM built upon a three-layer Boltzmann machine with Potts hidden units, which represent data clusters and classes. Through a statistical mechanics analysis, we derive saddle-point equations that characterize both the stationary points of DAMs trained on real data and the fixed points of DAMs trained on synthetic data within a teacher-student framework. Based on these results, we propose a novel regularization scheme that makes training significantly more stable. Moreover, we show empirically that our DAM learns interpretable solutions to both supervised and unsupervised classification problems. Pushing our theoretical analysis further, we find that the weights learned by relatively small DAMs correspond to unstable saddle points in larger DAMs. We implement a network-growing algorithm that leverages this saddle-point hierarchy to drastically reduce the computational cost of training dense associative memory.
@@ 1 Introduction
Studying the stationary points of machine learning algorithms is crucial to understand how they work. For example, [1, 2] demonstrated that local minima in the loss landscape of large artificial neural networks (NNs) are relatively close to the global minimum, explaining why they generalize well in practice. Moreover, [3, 4] showed that saddle points are much more numerous than local minima in large NNs. These breakthroughs were made by establishing deep connections between the loss landscape of NNs and the energy landscape of disordered systems studied in statistical mechanics. Beyond the broad insights provided by these studies, and despite the progress made by [5, 6, 7, 8], the classification of stationary points in machine learning algorithms remains an open problem. An interesting way forward is through dense associative memory (DAM) models [9, 10]. In fact, the interpretability of DAMs and the relative simplicity of their learning dynamics [11] suggest that their critical points could potentially be characterized using analytical calculations.
@@
The Hopfield network was originally introduced as a paradigmatic model of biological associative memory [9]. Generalized Hopfield networks [12, 13, 14, 15, 16, 17] were then developed to improve upon the limited storage capacity of the original [9, 18]. A few years ago, these generalized networks, commonly referred to as dense associative memory (DAM) or modern Hopfield networks [10, 19], were made into trainable machine learning models capable of accurate pattern classification by Krotov and Hopfield (K & H) [10]. In a nutshell, K & H’s DAM learns prototypes of patterns in a trainable weight matrix. Each prototype casts a vote for a class, and the patterns awaiting classification are assigned based on the votes of the prototypes that most closely resemble them. The resulting classification scheme is considerably more adversarially robust and interpretable than that of feedforward NNs with ReLU activation functions [10, 20, 21]. Since their debut as trainable machine learning architectures, deep connections have been made between modern Hopfield networks and transformers [19], as well as generative diffusion models [22, 23]. In particular, modern Hopfield networks were used to implement the attention mechanism of transformers [19, 24], which has been attracting a lot of interest in fundamental [25, 26, 27, 28, 29] and applied research [24, 28, 30, 31]. Recently, it was observed that the trainable weights of K & H’s DAM are channeled toward minima by a low-dimensional network of valleys in the loss landscape [11]. Moreover, the points where valleys branch out from one another were identified as saddles in the simple case where the DAM has two patterns to learn. In general, it is not straightforward to classify the stationary points of machine learning algorithms [8]. However, the results of [11] and the interpretability of DAMs suggest that their stationary points are both fundamental to their learning dynamics and easier to characterize than that of generic NNs. With this goal in mind, we revisit dense associative memory for pattern classification [10] using the framework of Boltzmann machines (BMs) [32, 33, 34, 35] and Statistical Mechanics.
In Section 2, we present the DAM model that we study and the tools that we use to do so. This Section is divided into two subsections. In Section 2.1, we derive a DAM from a BM template. In Section 2.2, we describe the setting in which we study DAM stationary points. In Section 3, we present our theoretical results. This Section is also divided into two subsections. In Section 3.1, we introduce saddle-point equations that characterize DAM stationary points. In Section 3.2, we use these equations to demonstrate the existence of a saddle-point hierarchy according to which the weights learned by DAMs of a given width are embedded in larger DAMs, where they become saddle points. In Section 4, we use our theoretical results to significantly improve DAM training, and we investigate the results. This Section is divided into three subsections. In Section 4.1, we design a theoretically motivated regularization method that facilitates supervised training. In Section 4.2, we show that our DAM, despite being designed for supervised learning, can learn interpretable solutions to both supervised and unsupervised classification problems. Finally, in Section 4.3, we relate our findings to the learning dynamics guided by valleys and saddles studied in [11], and we implement a network-growing algorithm [36, 37] that uses the saddle-point hierarchy to significantly reduce the cost of DAM training. The code and hyperparameter values used in our numerical experiments are available at this public repository [38].
@@ 2 Model
The Boltzmann machine (BM) is a canonical graphical model of correlations in discrete data [32]. It is customary to partition BMs into a visible layer $\mathbf{v}=\left\{v_{i}\right\}_{i=1}^{N}$ and a hidden layer $\mathbf{h}=\left\{h_{\mu}\right\}_{\mu=1}^{P}$ such that connections between the two layers are allowed, but connections within them are prohibited [33]. In this case, the visible layer represents concrete features of the data, whose mutual correlations are encoded in connections with the hidden layer. The restricted Boltzmann machine (RBM) obtained using this partition is much easier to train than a generic BM [34, 35] and still has considerable generating power [34, 39], making it more practical in machine learning applications [40, 41, 42]. The visible and hidden units of an RBM follow the Gibbs distribution
@@ 2.1 A dense associative memory (DAM) model
As mentioned in the Introduction, we will now derive a dense associative memory (DAM) model for classification from a BM template. We will explain why our model is a DAM at the end of this Section, once we have clearly defined it. We make three basic assumptions on the distribution of data to be classified:
@@
    the data is scale invariant, i.e. any data point $\mathbf{x}$ is equivalent to $c\mathbf{x}$;
@@
Since the data is scale invariant (Assumption 1), we normalize it by its (Euclidean) norm in the data layer. In other terms, we take the data units $x_{i}$ to be continuous variables with unit norm $\sqrt{\sum_{i=1}^{N}\left(x_{i}\right)^{2}}=1$. We assume no further knowledge about $\mathbf{x}$, so we take the prior $\operatorname{P}_{0}\left(\mathbf{x}\right)$ to be the uniform distribution on the N1N-1 dimensional unit sphere $S^{N-1}$. Data normalization is a very common practice in machine learning. For example, normalization by the Euclidean norm is frequent in text document clustering [44, 45, 46]. Various types of normalization also occur in the brain and retina [47].
Since the hidden layer and the class layer represent disjoint clusters and classes, respectively (Assumptions 2 and 3), we take their respective units to be mutually exclusive binary variables, i.e. $\mathbf{h}\in\{0,1\}^{P}$ with $\sum_{\mu=1}^{P}h_{\mu}\in\{0,1\}$ and $\mathbf{q}\in\{0,1\}^{C}$ with $\sum_{y=1}^{C}q_{y}\in\{0,1\}$. In other words, we take each of these two layers to be the vector representation of a single categorical (or Potts [48, 49]) variable with $P+1$ and $C+1$ categories, respectively. As such, $\operatorname{P}_{0}\left(\mathbf{h}\right)$ and $\operatorname{P}_{0}\left(\mathbf{q}\right)$ simplify to probability mass functions $\operatorname{P}_{0}\left(\mathbf{h}=\mathbf{e}_{\gamma}\right)$ and $\operatorname{P}_{0}\left(\mathbf{q}=\mathbf{e}_{y}\right)$, where we introduce $\mathbf{e}_{\gamma}=\left\{\delta_{\gamma\mu}\right\}_{\mu=1}^{P}$ for $\gamma\in\left\{0,\ldots,P\right\}$ and define $\mathbf{e}_{y}\in\{0,1\}^{C}$ analogously for $y\in\{0,\ldots,C\}$. In particular $\mathbf{e}_{0}=\mathbf{0}$ represents a state outside the $P$ clusters or the $C$ classes.
These priors on the hidden layer and the class layer can also be obtained by introducing fixed inhibitory connections within the hidden layer and the class layer, respectively [50, 51, 52]. Since at most one hidden unit $h_{\mu}$ can be activated at once, the hidden layer is a very sparse representation of the visible layer. In machine learning, sparsity can improve interpretability [53], generalization, computational efficiency [54], and adversarial robustness [55, 56, 57, 58, 59, 60]. The sparsity of the brain suggests that it is also beneficial for biological neural networks [61].
@@
In other words, the probability density $\operatorname{P}_{\beta}\left(\mathbf{x}|\mu,\mathbf{J}\right)$ corresponding to each cluster $\mu>0$ is a von Mises-Fisher (vMF) distribution centered on the direction $\mathbf{w}^{\mu}=\left\{w^{\mu}_{i}\right\}_{i=1}^{N}$ (see Appendix C). In order to interpret the $\mathbf{w}^{\mu}$ as centroids for their respective clusters, we assume that they belong to $S^{N-1}$ like the data layer. Under this assumption, we find the normalization constant of $\operatorname{P}_{\beta}\left(\mathbf{x}|\mu,\mathbf{J}\right)$ to be $\Omega_{N}\left(\beta\right)=\frac{\left(2\pi\right)^{N/2}I_{N/2-1}\left(\beta\right)}{\beta^{N/2-1}}$, where $I_{n}\left(x\right)$ is the modified Bessel function of the first kind of order $n$ (see Appendix C).
@@
where $\mathbf{p}=\left\{p^{\gamma}_{y}\right\}_{0\leq y\leq C}^{0\leq\gamma\leq P}=\left\{\operatorname{P}_{\beta}\left(\mathbf{q}=\mathbf{e}_{y},\mathbf{h}=\mathbf{e}_{\gamma}|\mathbf{J}\right)\right\}_{0\leq y\leq C}^{0\leq\gamma\leq P}$. $\Omega_{N}\left(0\right)=\frac{2\pi^{N/2}}{\Gamma\left(N/2\right)}$ is the surface area of $S^{N-1}$, where $\Gamma\left(x\right)$ is the Gamma function, so $\frac{1}{\Omega_{N}\left(0\right)}$ is the uniform distribution on $S^{N-1}$. The uniform distribution term of Eq. (2.1) encourages the model to ignore very noisy data during training, which may or not be desirable depending on the application.
The detailed derivation of Eq. (2.1) (in Appendix A) is inspired by the derivation of Gaussian mixtures from RBMs presented in [62] based on the framework of [50, 51, 52]. In our case, the marginal distribution $\operatorname{P}_{\beta}\left(\mathbf{x}|\mathbf{w},\mathbf{p}\right)=\sum_{y=0}^{C}\operatorname{P}_{\beta}\left(\mathbf{x},y|\mathbf{w},\mathbf{p}\right)$ and conditional distributions $\operatorname{P}_{\beta}\left(\mathbf{x}|y;\mathbf{w},\mathbf{p}\right)=\frac{\operatorname{P}_{\beta}\left(\mathbf{x},y|\mathbf{w},\mathbf{p}\right)}{\sum_{\gamma=0}^{P}\operatorname{P}_{\beta}\left(y,\gamma|\mathbf{w},\mathbf{p}\right)}$ are convex mixtures between the uniform distribution on $S^{N-1}$ and the vMF distributions $\operatorname{P}_{\beta}\left(\mathbf{x}|\mu,\mathbf{J}\right)$ (see Eq. 2). Probabilistic models similar to $\operatorname{P}_{\beta}\left(\mathbf{x}|\mathbf{w},\mathbf{p}\right)$ are notably used in text document clustering [46]. Mixture distributions that have class-dependent weights like $\operatorname{P}_{\beta}\left(\mathbf{x}|y;\mathbf{w},\mathbf{p}\right)$ are also used in Gaussian mixture discriminant analysis [63].
@@
which is a form of maximum likelihood estimation. We do so using constrained stochastic gradient descent with momentum (simply called SGD in this paper). We explain how the constraints on $\mathbf{w}$ and $\mathbf{p}$ are enforced in Appendix F, and we briefly discuss the initial conditions and the learning rate in Appendix I.
@@
High-dimensional probabilistic models that store or learn prototypes, such as our model, can typically reconstruct a limited number of patterns with non-trivial accuracy. In other words, they have limited storage capacity. For instance, Hopfield’s model of associative memory, the Hopfield network [9], has a capacity of up to $\mathcal{O}\left(N\right)$ patterns [9, 18]. Dense associative memory (DAM) models are a class of Hopfield network-inspired models with an asymptotically much higher capacity [10, 15]. The effective energy $\log\operatorname{P}_{\beta}\left(\mathbf{x}|y;\mathbf{w},\mathbf{p}\right)$ that we can minimize to reconstruct patterns with our model (see Eq. 2.1) is very similar to that of the DAM studied in [19, 27, 29], which also reconstructs patterns by energy minimization. In fact, according to [29], our model belongs to the class of DAMs with exponential capacity.

### v2

@@ Abstract
Dense associative memory (DAM) models have been attracting renewed attention since they were shown to be robust to adversarial examples and closely related to cutting edge machine learning paradigms, such as the attention mechanism and generative diffusion. We study a DAM built upon a three-layer Boltzmann machine with Potts hidden units, which represent data clusters and classes. Through a statistical mechanics analysis, we derive saddle-point equations that characterize both the stationary points of DAMs trained on real data and the fixed points of DAMs trained on synthetic data within a teacher-student framework. Based on these results, we propose a novel regularization scheme that makes training significantly more stable. Moreover, we show empirically that our DAM learns interpretable solutions to both supervised and unsupervised classification problems. Pushing our theoretical analysis further, we find that the weights learned by relatively small DAMs correspond to unstable saddle points in larger DAMs. We implement a network-growing algorithm that leverages this saddle-point hierarchy to drastically reduce the computational cost of training dense associative memory.
@@ 1 Introduction
Studying the stationary points of machine learning algorithms is crucial to understand how they work. For example, [1, 2] demonstrated that local minima in the loss landscape of large artificial Neural Networks (NNs) are relatively close to the global minimum, explaining why they generalize well in practice. Moreover, [3, 4] showed that saddle points are much more numerous than local minima in large NNs. These breakthroughs were made by establishing deep connections between the loss landscape of NNs and the energy landscape of disordered systems studied in statistical mechanics. Beyond the broad insights provided by these studies, and despite the progress made by [5, 6, 7, 8], the classification of stationary points in machine learning algorithms remains an open problem.
@@
Energy-based models have been playing a central role in studies of NNs and their theoretical properties. The Hopfield network, one of the most historically important energy-based models, was originally introduced as a paradigmatic model of biological associative memory [9]. Generalized Hopfield networks [10, 11] were then developed to improve upon the limited storage capacity of the original [9, 12]. The scale of the improvement was determined rigorously in following studies [13, 14, 15]. A few years ago, these generalized networks, commonly referred to as Dense Associative Memory (DAM) or modern Hopfield networks, were made into trainable machine learning models capable of accurate pattern classification by Krotov and Hopfield (K & H) [16]. In a nutshell, K & H’s DAM learns prototypes of patterns in a trainable weight matrix. Each prototype casts a vote for a class, and the patterns awaiting classification are assigned based on the votes of the prototypes that most closely resemble them. The resulting classification scheme is considerably more adversarially robust [17, 18] and interpretable [16] than that of feedforward NNs with ReLU activation functions. Since their debut as trainable machine learning architectures, deep connections have been made between modern Hopfield networks and other well-known and established machine learning paradigms, such as attention [19] and generative diffusion [20, 21]. In particular, modern Hopfield networks were used to implement the attention mechanism of transformers [19], which has been attracting a lot of interest in fundamental and applied research. See [22, 23] for a review of recent advances on the topic. Recently, it was observed that the trainable weights of K & H’s DAM are channeled toward minima by a low-dimensional network of valleys in the loss landscape [24]. Moreover, the points where valleys branch out from one another were identified as saddles in the simple case where the DAM has two patterns to learn. In general, it is not straightforward to classify the stationary points of machine learning algorithms [8]. However, the results of [24] and the interpretability of DAMs suggest that their stationary points are both fundamental to their learning dynamics and easier to characterize than that of generic NNs. With this goal in mind, we revisit dense associative memory for pattern classification [16] using the framework of Boltzmann Machines (BMs) [25, 26, 27, 28] and statistical mechanics.
The body of this paper is structured as follows. In Section 2, we introduce the DAM model that we study and the analytical tools that we use throughout our work. In particular, in Section 2.1, we derive the DAM in question from a BM template, and, in Section 2.2, we explain the setting that we use to analyze DAM stationary points.
Section 3 presents our theoretical contributions. In Section 3.1, we formulate saddle-point equations that characterize DAM stationary points, and, in Section 3.2, we leverage these equations to establish a saddle-point hierarchy principle stating that weights learned by DAMs of a given width are embedded as saddle points in wider DAMs.
Section 4 explains how the theoretical insights of Section 3 can be used to enhance training and explores how DAMs represent data once trained. In Section 4.1, we introduce a regularization method that facilitates supervised learning. Next, in Section 4.2, we show that our DAM, although designed for supervised tasks, can discover interpretable solutions in both supervised and unsupervised classification settings. Finally, in Section 4.3, we link our findings to the learning dynamics shaped by valleys and saddles studied in [24], and we implement a network-growing algorithm [29] that exploits the saddle-point hierarchy to significantly reduce DAM training costs.
The code and hyperparameter configurations used in our experiments are available in the following public repository [30].
@@ 2 Model
A Boltzmann Machine is a canonical graphical model of correlations in discrete data [25]. It is customary to partition BMs into a visible layer $\mathbf{v}=\left\{v_{i}\right\}_{i=1}^{N}\in\mathbb{R}^{N}$ and a hidden layer $\mathbf{h}=\left\{h_{\mu}\right\}_{\mu=1}^{P}\in\mathbb{R}^{P}$ such that connections between the two layers are allowed, but connections within them are prohibited [26]. In this case, the visible layer represents concrete features of the data, whose mutual correlations are encoded in connections with the hidden layer. The Restricted Boltzmann Machine (RBM) obtained using this partition is much easier to train than a generic BM [27, 28] and still has considerable generating power [27, 31], making it more practical in machine learning applications [32, 33, 34]. The visible and hidden units of an RBM follow the Gibbs distribution
@@ 2.1 A Dense Associative Memory (DAM) model
As mentioned in the Introduction, we will now derive a DAM model for classification from a BM template. We will explain why our model is a DAM at the end of this Section, once we have clearly defined it. We make three basic assumptions on the distribution of data to be classified:
@@
    the data is scale invariant, i.e. for all positive scalars $c$, data points $\mathbf{x}$ and $c\mathbf{x}$ are equivalent;
@@
Since the data is scale invariant (Assumption 1), the scale of individual data points contains no information about the classification task, so we normalize them by their (Euclidean) norm in the data layer. In other terms, we take the data units $x_{i}$ to be continuous variables with unit norm $\sqrt{\sum_{i=1}^{N}\left(x_{i}\right)^{2}}=1$. We assume no further knowledge about $\mathbf{x}$, so we take the prior $\operatorname{P}_{0}\left(\mathbf{x}\right)$ to be the uniform distribution on the N1N-1 dimensional unit hypersphere $S^{N-1}$, i.e. is the set of all $\mathbf{x}$ with $\sqrt{\sum_{i=1}^{N}\left(x_{i}\right)^{2}}=1$. Data normalization is a very common practice in machine learning. For example, normalization by the Euclidean norm has been popular in text document clustering even since its introduction in the 1980s [36]. Various types of normalization also occur in the brain and retina [37].
Since the hidden layer and the class layer aim to represent disjoint clusters and classes, respectively (Assumptions 2 and 3), we take their respective units to be mutually exclusive binary variables, i.e. $\mathbf{h}\in\{0,1\}^{P}$ with $\sum_{\mu=1}^{P}h_{\mu}\in\{0,1\}$ and $\mathbf{q}\in\{0,1\}^{C}$ with $\sum_{y=1}^{C}q_{y}\in\{0,1\}$. At any given time, at most one unit per layer can take the value $1$, representing the fact that the clusters and classes are disjoint. In other words, we take each of the two layers to be the vector representation of a single categorical (or Potts [38, 39]) variable with $P+1$ and $C+1$ categories, respectively. As such, $\operatorname{P}_{0}\left(\mathbf{h}\right)$ and $\operatorname{P}_{0}\left(\mathbf{q}\right)$ simplify to probability mass functions $\operatorname{P}_{0}\left(\mathbf{h}=\mathbf{e}_{\gamma}\right)$ and $\operatorname{P}_{0}\left(\mathbf{q}=\mathbf{e}_{y}\right)$, where we introduce $\mathbf{e}_{\gamma}=\left\{\delta_{\gamma\mu}\right\}_{\mu=1}^{P}$ for $\gamma\in\left\{0,\ldots,P\right\}$ and define $\mathbf{e}_{y}\in\{0,1\}^{C}$ analogously for $y\in\{0,\ldots,C\}$. In particular $\mathbf{e}_{0}=\mathbf{0}$ represents a state outside the $P$ clusters or the $C$ classes.
These priors on the hidden layer and the class layer can also be obtained by introducing fixed inhibitory connections within the hidden layer and the class layer, respectively [40, 41]. Since at most one hidden unit $h_{\mu}$ can be activated at once, the hidden layer is a very sparse representation of the visible layer. In machine learning, sparsity can improve interpretability [42], generalization, computational efficiency [43], and adversarial robustness [44]. The sparsity of the brain suggests that it is also beneficial for biological neural networks [45].
@@
In other words, the probability density $\operatorname{P}_{\beta}\left(\mathbf{x}|\mu,\mathbf{J}\right)$ corresponding to each cluster $\mu>0$ is a von Mises-Fisher (vMF) distribution centered on the direction $\mathbf{w}^{\mu}=\left\{w^{\mu}_{i}\right\}_{i=1}^{N}$ (see Appendix B). In order to interpret the $\mathbf{w}^{\mu}$ as centroids for their respective clusters, we assume that they belong to $S^{N-1}$ like the data layer. Under this assumption, we find the normalization constant of $\operatorname{P}_{\beta}\left(\mathbf{x}|\mu,\mathbf{J}\right)$ to be $\Omega_{N}\left(\beta\right)=\frac{\left(2\pi\right)^{N/2}I_{N/2-1}\left(\beta\right)}{\beta^{N/2-1}}$, where $I_{n}\left(x\right)$ is the modified Bessel function of the first kind of order $n$ (see Appendix B).
@@
where $\mathbf{p}=\left\{p^{\gamma}_{y}\right\}_{0\leq y\leq C}^{0\leq\gamma\leq P}=\left\{\operatorname{P}_{\beta}\left(\mathbf{q}=\mathbf{e}_{y},\mathbf{h}=\mathbf{e}_{\gamma}|\mathbf{J}\right)\right\}_{0\leq y\leq C}^{0\leq\gamma\leq P}$. $\Omega_{N}\left(0\right)=\frac{2\pi^{N/2}}{\Gamma\left(N/2\right)}$ is the surface area of $S^{N-1}$, where $\Gamma\left(x\right)$ is the Gamma function, so $\frac{1}{\Omega_{N}\left(0\right)}$ is the uniform distribution on $S^{N-1}$. The uniform distribution term of Eq. (2.1) encourages the model to ignore very noisy data during training, which may or may not be desirable depending on the application.
The detailed derivation of Eq. (2.1) (in Appendix A) is inspired by the derivation of Gaussian mixtures from RBMs presented in [46] based on the framework of [40, 41]. In our case, the marginal distribution $\operatorname{P}_{\beta}\left(\mathbf{x}|\mathbf{w},\mathbf{p}\right)=\sum_{y=0}^{C}\operatorname{P}_{\beta}\left(\mathbf{x},y|\mathbf{w},\mathbf{p}\right)$ and conditional distributions $\operatorname{P}_{\beta}\left(\mathbf{x}|y;\mathbf{w},\mathbf{p}\right)=\frac{\operatorname{P}_{\beta}\left(\mathbf{x},y|\mathbf{w},\mathbf{p}\right)}{\sum_{\gamma=0}^{P}\operatorname{P}_{\beta}\left(y,\gamma|\mathbf{w},\mathbf{p}\right)}$ are convex mixtures between the uniform distribution on $S^{N-1}$ and the vMF distributions $\operatorname{P}_{\beta}\left(\mathbf{x}|\mu,\mathbf{J}\right)$ (see Eq. 2). Probabilistic models similar to $\operatorname{P}_{\beta}\left(\mathbf{x}|\mathbf{w},\mathbf{p}\right)$ are notably used in text document clustering [47]. Mixture distributions that have class-dependent weights like $\operatorname{P}_{\beta}\left(\mathbf{x}|y;\mathbf{w},\mathbf{p}\right)$ are also used in Gaussian mixture discriminant analysis [48].
@@
which is a form of maximum likelihood estimation. We do so using constrained Stochastic Gradient Descent with momentum (simply called SGD in this paper). We explain how the constraints on $\mathbf{w}$ and $\mathbf{p}$ are enforced in Appendix C, and we briefly discuss the initial conditions and the learning rate in Appendix D.
@@
High-dimensional probabilistic models that store or learn prototypes, such as our model, can typically reconstruct a limited number of patterns with nontrivial accuracy. In other words, they have limited storage capacity. For instance, Hopfield’s model of associative memory, the Hopfield network [9], has a capacity of up to $\mathcal{O}\left(N\right)$ patterns [9, 12]. Dense Associative Memory (DAM) models, which are inspired by the Hopfield network, are a class of models with an asymptotically much higher capacity [14, 16]. The effective energy $\log\operatorname{P}_{\beta}\left(\mathbf{x}|y;\mathbf{w},\mathbf{p}\right)$ that we can minimize to reconstruct patterns with our model (see Eq. 2.1) is very similar to that of the DAM with exponential capacity introduced in [19], which uses energy minimization to implement the attention mechanism of transformers. In fact, according to [51], our model also belongs to the class of DAMs with exponential capacity.